2. Find a model for each of the following wffs. a. 3x (p(x) — 9(x))^VX - p(x) b. 3x Vy p(x, y) ^ 3x Vy - p(x, y)

Need help please
Answer #1

Answers @ 3x (PCW) ->9(29)1trx -D COD let the domain be D-{a,b} p(x) and qa) be such that let pla) =E, Pb) = F gea) -T, 9(0)=p(b,a) = F , pcb, b) = F. Then, for x-a, Hy pex, y) [since pla,al=T= p (a, b)] Also, for 2 =b, Ay - pcx, y) [since plbia)=F=p